# Guided Interpretable Facial Expression Recognition via Spatial Action Unit Cues

### v1

@@ 2 Related Work
Action units in FER systems. The Facial Action Coding System (FACS) is a taxonomy for fine-grained facial expression analysis [22, 27]. They have long been used to analyze facial expressions [23, 32, 48, 65, 91]. Each basic facial expression is associated with a list of action units as they determine which facial muscles are involved in expressing such emotion. The common established task in the literature is action units detection [32, 48]. It is a supervised multi-label classification task. It aims at predicting the correct set of active action units in an input image. To perform such a task, a tedious annotation is required to determine the active set of action units in each image since not all action units must be active at once. Another related task goes further to estimate the intensity of the action units [23, 91] which is furthermore challenging. Other works aim to jointly localize and estimate the intensity [65]. The work of [32] is relatively close to our work. In order to accurately detect active action units through multi-label classification, authors leverage multi-task learning to jointly predict a localization heatmap for the same action units. The goal is to improve action units detection via their spatial localization. However, action units have not been used for the interpretability of FER classifiers. This makes our work the first to do so. To avoid the extra annotation cost burden, we use generic discriminative action units codebook defined in [51]. It allows the automatic labeling of large benchmarks without manual intervention (Fig. 3). The codebook associates a set of action units to a basic facial expression (Fig. 2). This is used to build a discriminative heatmap that holds potential ROIs to be inspected by the model in order to determine the facial expression in the input image. This does not require all the action units to be active. But, it is more likely that part of them are active in order for the expression to be manifested in an image. A further step of our work is to use active action units to build an ideal interpretable model but it comes at an additional expensive annotation cost. Alternatively, a pretrained action units detector could be used such as [32].
@@
Figure 3: Our interpretable classifier for the FER task (training and inference). Each basic facial expression can be determined via a set of action units [51]. Therefore, to train our interpretable classifier, we first extract facial landmarks and build a discriminative spatial map ${\bm{A}}$ that contains the set of all action units associated with the image class expression [51]. This map is used as localization cues to train layer-wise attention ${\bm{T}_{l}}$ to focus on the ROIs highlighted in the action units map. Classification loss, such as cross-entropy, is also used. Once trained, the classifier yields an interpretable layer-wise attention map. When CAM-based classifier [62] is considered, the model can also produce a per-class interpretable map (i.e., CAM).
@@ 3.1 Notation
3.2 Building spatial action units maps
Given an input image and its label ${(\bm{X},y)}$, standard 68 facial landmarks are extracted from the image using off-the-shelf techniques such as SynergyNet [76]. This allows us to localize relevant parts of the face such as mouth, nose, eyes, and eyebrows, which will be helpful later to estimate the location of relevant action units. Martinez et al. [51] suggest that each facial expression can be recognized by inspecting a combination of action units (Fig. 2). For instance, the expression ’Happiness’ involves a set of action units: ’cheek raise’, ’lip corner puller’, ’lips part’. Using the extracted facial landmarks around the mouth part, we can localize the ’lip part’ of the face. Such localization is translated into a 2D heatmap where strong activations are concentrated around the ’lip’ to indicate an ROI, and tiny activation everywhere else to indicate background. Such a 2D map is of great importance in training since it provides useful localization cues for relevant discriminative parts in an input image for the FER task used by experts. Typically, this requires manual annotation to delineate ROIs, increasing the annotation cost. Unfortunately, such valuable annotation is not available in public facial expression benchmarks. This work simulates such ROIs without an additional manual supervision cost. Using the codebook of facial expressions and their associated action units provided in [51] (Fig. 2), in combination with the extracted facial landmarks and the image label expression ${y}$, a single 2D heatmap, ${\bm{A}\in\mathbb{R}^{h\times w}}$, is built to hold all the relevant (i.e., discriminative) action units for the input image (Fig. 3). Consequently, the input image is augmented with an extra supervision cue to be ${(\bm{X},y,\bm{A})}$ where ${\bm{A}}$ is the estimated action units map. To ease subsequent alignment computations, this map is normalized to have values between 0 and 1: ${\bm{A}\in[0,1]^{h\times w}}$. Note that this process requires only the image class, i.e. image facial expression, as supervision. No extra manual annotations are required.
@@ 3.3 Layer-wise attention
where $n$ is the total number of feature maps. The discrepancy in activations in ${\bm{T}_{l}}$ differentiates relevant from irrelevant regions. High activations indicate potential spatial ROIs for classification, while low activations point to background and noisy regions. In a deep classifier, such attention maps are a great candidate to incorporate spatial learning cues into the model. We note that although CAM tensors ${\bm{M}}$ hold spatial discriminative cues, they are very sensitive to being altered. Aligning them explicitly with other maps can easily lead to poor classification [7, 12]. Therefore, we rely on layer-wise attention to learn better and interpretable spatial features in this work.
3.4 Action units alignment loss
After estimating both layer-wise attention ${\bm{T}_{l}}$ and action units map ${\bm{A}}$, we can proceed with their alignment to train the classifier to yield similar spatial cues. Since the dynamic range of the attention map is unknown beforehand, and since it changes during training, it is inadequate to train the attention to have the same values as the action unit map ${\bm{A}}$. Instead, we resort to a loss that aims to yield attention maps that are correlated with ${\bm{A}}$. To this end, we use cosine similarity as,
@@ 3.5 Total training loss
Given the triplet ${(\bm{X},y,\bm{A})}$, our goal is to train the classifier ${f}$ to correctly classify the input image ${\bm{X}}$ to yield the right class ${y}$. Additionally, we aim to push the layer ${l}$ to construct spatial features that are correlated with the action units map cues ${\bm{A}}$. To achieve this, we jointly minimize the following composite loss,
@@
where ${\lambda\geq 0}$ is a weighting coefficient that balances the importance of attention alignment with the action units map compared to cross-entropy loss (left side term). The generalization of Eq. 3 to a combination of multiple layers is straightforward. It can be achieved by adding more layer-wise terms to the loss. Minimizing Eq. 3 incorporates explicitly the experts’ procedure in assessing basic facial expressions in images into the model decision process. The classifier is trained to localize the relevant action units in order to build discriminative features for expression prediction. This justifies using layer-wise attention ${\bm{T}_{l}}$ as an interpretability tool during inference time (Fig. 3).
We note that the computational cost added by the alignment term is negligible since it can be easily computed on GPU. This leads to a training time similar to the vanilla case where no alignment is used. The only required pre-processing is facial landmarks extraction which can be done off-line once and stored on disk. Action units map ${\bm{A}}$ can be computed in a negligible time on CPU on the fly during training.
@@
Table 1: Classification (CL) and CAM-localization (CAM-COS) performance over RAF-DB and AffectNet test sets with and without action units (AUs) across different methods.
@@
Table 2: Attention-localization (ATT-COS) (at layer 5) performance over RAF-DB and AffectNet test sets with and without action units (AUs).
4 Results and Discussion
4.1 Experimental Methodology
- RAF-DB Real-world Affective Faces Database is a large scale facial expression dataset [39]. All images have been manually annotated by multiple annotators. It contains six basic expressions (’Happiness’, ’Sadness’, ’Surprise’, ’Anger’, ’Disgust’, ’Fear’) and ’Neutral’. The dataset contains 12,271 samples for training and 3,068 samples for testing. Both sets have the same distribution.
- AffectNet is one of the largest facial expression recognition datasets with 420k images were manually annotated. Following [37], 283,901 images are used for training, and 3,500 images for testing. In terms of labels, this dataset has the same six basic expressions as in the RAF-DB dataset, in addition to ’Neutral’.
Implementation Details. For all our experiments, we trained each method for 100 epochs over RAF-DB with a mini-batch size of 32 and 20 epochs for AffectNet, with a mini-batch size of 224. For CNN-based methods, we used ResNet50 [31] as an encoder. For transformer-based methods, we used Deit-S [70] for TS-CAM [29]. For the APVIT method [82], we used their proposed transformer architecture which also relies on Deit-S [70]. Images are aligned and resized to ${256\times 256}$, and randomly cropped patches of size ${224\times 224}$ are extracted for training. For APVIT [82], images are resized to ${112\times 112}$ and fed as input following the method’s protocol. The hyper-parameter ${\lambda>0}$ is estimated using validation by covering values that range from less than 1 to 20. The optimization of training loss is performed using stochastic gradient descent (SGD). The hyper-parameters search of different methods is done over the RAF-DB dataset. Due to its large size, relevant hyper-parameters are transferred from RAF-DB to AffectNet dataset.
Baseline Methods. To assess the benefits of using action unit cues, we leverage WSOL classifiers. They allow the classification of an image using only the image class as a label. In addition, they provide interpretability maps. In particular, we use CAM-based WSOL methods, which provide interpretability by highlighting ROIs associated with a class. To this end, we selected methods that cover both families of CAM techniques [62]: 1- Bottom-up methods where information flows from the input layer to the output layer. This includes CAM [93], WILDCAT [20], ACoL [89], PRM [95], ADL [13], and TS-CAM [29]. 2- Top-down methods where information flows in both directions. This includes gradient-based methods: GradCAM [66], GradCAM++ [10], and LayerCAM [34]. All these methods are common in the WSOL field. In the FER task, we selected a recent state-of-the-art method, APVIT [82], which builds spatial attention in its architecture to yield interpretable maps. However, it does not have a CAM module. All the methods are trained with and without our AUs loss in addition to classification loss in order to assess the impact of using action units as spatial cues for learning. We note that all results are reported using our own implementations.
To build action units map cues, we first extract standard 68 facial landmarks from images using SynergyNet [76]. Following [51], each facial expression can be recognized using only a set of generic action units. Using the extracted facial landmarks, one can build a 2D map that contains a heatmap of a single action unit. For an image, a single 2D heatmap is then built containing all the action units associated with the image expression. This heatmap encodes the location of the necessary discriminative spatial ROIs needed to recognize the facial expression presented in the image. Training a classifier to focus mainly only on these spatial ROIs in an image is expected to improve its interpretability and, furthermore, its classification accuracy. Note that the only manual supervision required for our method is the image expression. No extra manual annotations are required.
Evaluation Metrics. We use standard classification accuracy (CL) for the classification task. We also report a classification confusion matrix. Localization of action units is performed at CAM, and layer-wise attention maps. Although they are both normalized between [0, 1], they do not necessarily have the same exact value at the pixel level, compared to action units map, since attention is trained to be correlated with action unit maps. Therefore, we resort to using cosine similarity (Eq. 2) [^2] For evaluating a map, either a CAM or an attention map, it is first upsampled to the same size as the action units map ${\bm{A}}$. as a localization (interpretability) measure to assess how a CAM or attention is well aligned with the action unit map. Higher similarity indicates better localization and, hence, better interpretability. We refer to the cosine similarity between a CAM and an action units map as CAM-COS and as ATT-COS for the case of attention map versus action units map. Similarly to the classification confusion matrix, we report a per-class cosine similarity matrix. We note that during training, we perform the model selection by taking the model with the best classification accuracy (CL), and we report its corresponding localization metrics.
4.2 Comparison with different methods
Tab.1 shows the impact of using AUs alignment over classification and CAM interpretability across different methods. Overall, using our proposed action units alignment helped improving the CAM interpretability with large margin. Since CAMs are often built after the last layer where we applied the alignment, they are strongly affected leading to better AUs localization. We note that classification accuracy (CL) has also improved with relatively large margin over AffectNet compared to RAF-DB. This may be related to the fact there is still a large room for improvement in AffectNet compared to RAF-DB. In term of attention interpretability (Tab.2), the effect of using AUs loss is quite larger than in CAM. Most importantly, since the alignment is applied over layer-wise attention, it led to an attention that is better in interpretability than CAM. We note that achieving better classification accuracy does not necessarily mean better interpretability. For instance, vanilla APVIT method [82] has the highest classification accuracy CL of ${91\%}$ over RAF-DB. However, it only scores 0.380.38{0.38}0.38 in term of attention interpretability ATT-COS, the lowest score across all models. Since vanilla case is optimized only to maximize classification score, it can easily find features, that are not necessarily interpretable, to achieve that. Interestingly, we show that the same model can achieve similar classification accuracy of ${91.03\%}$ and an attention interpretability score of 0.850.85{0.85}0.85 when using our alignment loss. This suggests that spatial features have largely shifted their focus, as illustrated in Fig. 6, without compromising classification performance. As a result, the model finds different spatial features that satisfies both losses: classification and interpretabilty (AUs). We provide in Fig.4 an illustration of prediction with and without our action units alignment. In addition, Figs. 6, 7, 8, 9 show per-class average attention and CAM over both dataset. These visuals show better improvement of our classifier interprebility compared to vanilla case. We further more report per-class classification accuracy and localization over both dataset in Fig.10, 11. Overall, and across both dataset, we observe that the class ’Disgust’ is the most difficult to localize its spatial cues.
Figure 5: Ablations on the RAF-DB test set: impact of ${\lambda}$ over classification and localization (intepretability) performance. Alignment to AUs is performed over layer 5 of ResNet50 [31] with CAM [93] method.
@@ 4.2 Comparison with different methods
Table 3: Ablation study over RAF-DB test set: impact of applying AUs alignment over different layers over classification (CL) and CAM-localization (CAM-COS). Method: CAM [93] and ResNet50 [31] encoder. Layer 1 is the input layer.

### v3

@@ 2.4 AUs in FER systems
The Facial Action Coding System (FACS) is a taxonomy for fine-grained facial expression analysis [22, 27]. They have long been used to analyze facial expressions [23, 32, 48, 65, 91]. Each basic facial expression is associated with a list of AUs as they determine which facial muscles are involved in expressing such emotion. The standard established task in the literature is AU detection [32, 48]. It is a supervised multi-label classification task. It aims at predicting the correct set of active AUs in an input image. To perform such a task, a tedious annotation is required to determine the active set of AUs in each image since not all AUs must be active at once. Another related task goes further to estimate the intensity of the AUs [23, 91], which is furthermore challenging. Other works aim to localize and estimate the intensity [65] jointly. The work of [32] is relatively close to our work. To accurately detect active AUs through multi-label classification, authors leverage multi-task learning to predict a localization heatmap for the same AUs jointly. The goal is to improve the detection of AUs via their spatial localization. However, AUs have not been used for the interpretability of FER classifiers. This makes our work the first to do so. To avoid the extra annotation costs, we used a generic discriminative AUs codebook defined in [51]. It allows the automatic labeling of extensive benchmarks without manual intervention (Fig. 3). The codebook associates a set of AUs to a basic facial expression (Fig. 2). This is used to build a discriminative heatmap that holds potential ROIs to be inspected by the model to determine the facial expression in the input image. This does not require all the AUs to be active. But, it is more likely that part of them are active for the expression to manifest in an image. A further step of our work is to use active AUs to build an ideal interpretable model, although this comes at an additional expensive annotation cost. Alternatively, a pretrained AU detector, such as [32], could be used.
@@
Figure 3: Our interpretable classifier for the FER task (training and inference). Each basic facial expression can be determined via a set of AUs [51]. Therefore, to train our interpretable FER classifier, we first extract facial landmarks and build a discriminative spatial map ${\bm{A}}$ that contains the set of all AUs associated with the image class expression [51]. This map is used as localization cues to train layer-wise attention ${\bm{T}_{l}}$ to focus on the ROIs highlighted in the AU map. A classification loss, such as cross-entropy, is also used. Once trained, the classifier yields an interpretable layer-wise attention map. When a CAM method [62] is considered, the classifier can also produce a per-class interpretable map.
@@ 3.1 Notation
3.2 Building spatial AU maps
Given an input image and its label ${(\bm{X},y)}$, standard 68 facial landmarks are extracted from the image using off-the-shelf techniques such as SynergyNet [76]. This allows us to localize relevant parts of the face such as mouth, nose, eyes, and eyebrows, which will be helpful later to estimate the location of relevant AUs. Martinez et al. [51] suggest that each facial expression can be recognized by inspecting a combination of AUs (Fig. 2). For instance, the expression ’Happiness’ involves a set of AUs: ’cheek raise’, ’lip corner puller’, ’lips part’. Using the extracted facial landmarks around the mouth part, we can localize the ’lip part’ of the face. Such localization is translated into a 2D heatmap where strong activations are concentrated around the ’lip’ to indicate an ROI and tiny activations everywhere else to indicate the background. Such a 2D map is of great importance in training since it provides useful localization cues for relevant discriminative parts in an input image for the FER task used by experts. Typically, this requires manual annotation to delineate ROIs, increasing the annotation cost. Unfortunately, such valuable annotation is not available in public facial expression benchmarks. This work simulates such ROIs without an additional manual supervision cost. Using the codebook of facial expressions and their associated AUs provided in [51] (Fig. 2), in combination with the extracted facial landmarks and the image label expression ${y}$, a single 2D heatmap, ${\bm{A}\in\mathbb{R}^{h\times w}}$, is built to hold all the relevant (i.e., discriminative) AUs for the input image (Fig. 3). Consequently, the input image is augmented with an extra supervision cue to be ${(\bm{X},y,\bm{A})}$ where ${\bm{A}}$ is the estimated AU map. To ease subsequent alignment computations, this map is normalized to have values between 0 and 1: ${\bm{A}\in[0,1]^{h\times w}}$. Note that this process requires only the image class, i.e., image facial expression, as supervision. No extra manual annotations are needed.
@@ 3.3 Layer-wise attention
where $n$ is the total number of feature maps. The discrepancy in activations in ${\bm{T}_{l}}$ differentiates relevant from irrelevant regions. High activations indicate potential spatial ROIs for classification, while low activations point to background and noisy regions. In a deep classifier, such attention maps are a great candidate for incorporating spatial learning cues into the model. Although CAM tensors ${\bm{M}}$ hold spatial discriminative cues, they are sensitive to being altered. Aligning them explicitly with other maps can easily lead to poor classification [7, 12]. Therefore, we rely on layer-wise attention to learn better and interpretable spatial features.
3.4 AU map alignment loss
After estimating both layer-wise attention ${\bm{T}_{l}}$ and AU map ${\bm{A}}$, we can proceed with their alignment to train the classifier to yield similar spatial cues. Since the dynamic range of the attention map is unknown beforehand, and since it changes during training, it is inadequate to train the attention to have the same values as the AU map ${\bm{A}}$. Instead, we resort to a loss that aims to yield attention maps that are correlated with ${\bm{A}}$. To this end, we use cosine similarity as,
@@ 3.5 Total training loss
Given the triplet ${(\bm{X},y,\bm{A})}$, our goal is to train the classifier ${f}$ on input image ${\bm{X}}$ to yield the correct class ${y}$. Additionally, we aim to encourage layer ${l}$ to construct spatial features that are correlated with the AU map cues ${\bm{A}}$. To achieve this, we jointly minimize the following composite loss,
@@
where ${\lambda\geq 0}$ is a weighting coefficient that balances the importance of attention alignment with the AU map compared to cross-entropy loss (left side term). The generalization of Eq. (3) to a combination of multiple layers is straightforward. It can be achieved by adding more layer-wise terms to the loss. Minimizing Eq. (3) incorporates explicitly the experts’ procedure in assessing basic facial expressions in images into the model decision process. The classifier is trained to localize the relevant AUs to build discriminative features for expression prediction. This justifies using layer-wise attention ${\bm{T}_{l}}$ as an interpretability tool during inference time (Fig. 3).
We note that the computational cost added by the alignment term is negligible since it can be easily computed on GPU. This leads to a training time similar to the vanilla case, where no alignment is used. The only required pre-processing is facial landmark extraction, which can be done offline once and stored on disk. AU map ${\bm{A}}$ can be computed in a negligible time on CPU on the fly during training using the image-class label only as supervision. Since image-class labels are unavailable at test time, AU maps can not be built using the lookup table emotion-AUs. Hence, training the model to produce them along with the true label is more realistic and practical. Furthermore, it allows the model itself to build them, leading to more interpretable predictions.
@@
Table 1: Classification (CL) and CAM-localization (CAM-COS) performance on RAF-DB and AffectNet test sets with and without AUs across methods.
@@
Table 2: Attention-localization (ATT-COS) (at layer 5) performance over RAF-DB and AffectNet test sets with and without AUs.
@@
Table 3: Ablation study over RAF-DB test set: impact of applying AU alignment over different layers over classification (CL) and CAM-localization (CAM-COS). Method: CAM [93] and ResNet50 [31] encoder. Layer 1 is the input layer.
4 Results and Discussion
4.1 Experimental Methodology
- RAF-DB Real-world Affective Faces Database is a large scale facial expression dataset [39]. Multiple annotators have manually annotated all images. It contains six basic expressions (’Happiness’, ’Sadness’, ’Surprise’, ’Anger’, ’Disgust’, ’Fear’) and ’Neutral’. The dataset contains 12,271 samples for training and 3,068 samples for testing. Both sets have the same distribution.
- AffectNet is one of the largest facial expression recognition datasets with 420k images were manually annotated. Following [37], 283,901 images are used for training and 3,500 for testing. In terms of labels, this dataset has the same six basic expressions as in the RAF-DB dataset, in addition to ’Neutral’.
Implementation Details. For all our experiments, we trained each method for 100 epochs over RAF-DB with a mini-batch size 32 and 20 epochs for AffectNet, with a mini-batch size 224. For CNN-based methods, we used ResNet50 [31] as an encoder. For transformer-based methods, we used Deit-S [70] for TS-CAM [29]. We used their proposed transformer architecture for the APVIT method [82], which also relies on Deit-S [70]. Images are aligned and resized to ${256\times 256}$, and randomly cropped patches of size ${224\times 224}$ are extracted for training. For APVIT [82], images are resized to ${112\times 112}$ and fed as input following the method’s protocol. The hyper-parameter ${\lambda>0}$ is estimated using validation by covering values that range from less than 1 to 20. The optimization of training loss is performed using stochastic gradient descent (SGD). The hyper-parameters search of different methods is done over the RAF-DB dataset. Due to its large size, relevant hyper-parameters are transferred from RAF-DB to AffectNet dataset.
Baseline Methods. To assess the benefits of using AU cues, we leverage WSOL classifiers. They allow the classification of an image using only the image class as a label. In addition, they provide interpretability maps. In particular, we use CAM-based WSOL methods, which provide interpretability by highlighting ROIs associated with a class. To this end, we selected methods that cover both families of CAM techniques [62]: 1- Bottom-up methods where information flows from the input layer to the output layer. This includes CAM [93], WILDCAT [20], ACoL [89], PRM [95], ADL [13], and TS-CAM [29]. 2- Top-down methods where information flows in both directions. This includes gradient-based methods: GradCAM [66], GradCAM++ [10], and LayerCAM [34]. All these methods are common in the WSOL field. In the FER task, we selected a recent state-of-the-art method, APVIT [82], which builds spatial attention in its architecture to yield interpretable maps. However, it does not have a CAM module. All the methods are trained with and without our AU loss in addition to classification loss to assess the impact of using AUs as spatial cues for learning. We note that all results are reported using our implementations.
To build AU map cues, we first extract standard 68 facial landmarks from images using SynergyNet [76] [^2] Other facial landmarks extractors can also easily be used used.. Following [51], each facial expression can be recognized using only a set of generic AUs. Using the extracted facial landmarks, one can build a 2D map that contains a heatmap of a single AU. For an image, a single 2D heatmap is then built containing all the AUs associated with the image expression. This heatmap encodes the location of the necessary discriminative spatial ROIs needed to recognize the facial expression presented in the image. Training a classifier to focus mainly on these spatial ROIs in an image is expected to improve its interpretability and classification accuracy. Note that the only manual supervision required for our method is the image expression. No extra manual annotations are required.
Evaluation Metrics. We use standard classification accuracy (CL) for the classification task. We also report a classification confusion matrix. Localization of AUs is performed at CAM, and layer-wise attention maps. Although they are both normalized between [0, 1], they do not necessarily have the same value at the pixel level, compared to AU map, since attention is trained to be correlated with AU maps. Therefore, we resort to using cosine similarity (Eq. 2) as a localization (interpretability) measure to assess how a CAM or attention is well aligned with the AU map. To evaluate either a CAM or an attention map, it is first upsampled to the same size as the AU map ${\bm{A}}$. Higher similarity indicates better localization and, hence, better interpretability. We refer to the cosine similarity between a CAM and an AU map as CAM-COS and as ATT-COS for the case of attention map versus AU map. Similarly to the classification confusion matrix, we report a per-class cosine similarity matrix. During training, we select the model with the best classification accuracy (CL), and report its corresponding localization metrics.
4.2 Comparison with different methods
Table 1 shows the impact of using AU alignment over classification and CAM interpretability across different methods. Overall, using our proposed AU alignment helped improve the CAM interpretability by a large margin. Since CAMs are often built after the last layer where we applied the alignment, they are strongly affected, leading to better AU localization. We note that classification accuracy (CL) has also improved with a relatively large margin over AffectNet compared to RAF-DB. This may be related to the fact there is still a large room for improvement in AffectNet compared to RAF-DB. Regarding attention interpretability (Table 2), the effect of using AU loss is larger than in CAM. Most importantly, since the alignment is applied over layer-wise attention, it leads to attention that is better in interpretability than CAM. Note that achieving better classification accuracy does not necessarily translate to better interpretability. For instance, the vanilla APVIT method [82] has the highest classification accuracy CL of ${91\%}$ over RAF-DB. However, it only scores 0.380.38{0.38}0.38 in terms of attention interpretability ATT-COS, the low est score across all models. Since the vanilla case is optimized only to maximize the classification score, it can easily find features that are not necessarily interpretable to achieve that. Interestingly, we show that the same model can achieve a similar classification accuracy of ${91.03\%}$ and an attention interpretability score of 0.850.85{0.85}0.85 when using our alignment loss. This suggests that spatial features have largely shifted their focus, as illustrated in Fig. 6, without compromising classification performance. As a result, the model finds different spatial features that satisfy both losses: classification and interpretability (AU). On RAF-DB and AffectNet, we note that our best obtained classification performance of ${91.03\%}$ and ${62.28\%}$ are close to the current state-of-the-art performance of ${92.21\%}$ and ${67.49\%}$ achieved by [50], respectively. Additionally, building AU maps during training adds negligible computational time. On the RAF-DB dataset, we obtain a training time of 1min 19sec /epoch and 1min 47sec /epoch without and with AU maps. Over the AffectNet dataset, rates of 15min 9sec /epoch and 17min 35sec /epoch are achieved without and with AU maps. We provide in Fig.4 an illustration of prediction with and without our action units alignment. In addition, Figs. 6, 7, 8, 9 show per-class average attention and CAM over both dataset. These visuals show better improvement of our classifier interprebility compared to vanilla case.
Figure 5: Ablations on the RAF-DB test set: impact of ${\lambda}$ over classification and localization (interpretability) performance. Alignment to AUs is performed over layer 5 of ResNet50 [31] with CAM [93] method.